Ivan and Kate live 40 miles apart. Ivan leaves his house at 8:00 a.m. and bicycles towards Kate's house at a constant speed of 12 mph. Kate leaves her house at 9:30 a.m. bicycles towards Ivan at constant speed of 16 mph. At what time will they meet?
Answer by stanbon(75887)

Ivan and Kate live 40 miles apart. Ivan leaves his house at 8:00 a.m. and bicycles towards Kate's house at a constant speed of 12 mph. Kate leaves her house at 9:30 a.m. bicycles towards Ivan at constant speed of 16 mph. At what time will they meet?
-----
Ivan DATA:
rate = 12 mph ; time = t hrs ; distance = 12t miles
--------
Kate DATA:
rate = 16 mph ; time = t-(3/2) hrs ; distance = 16t - 24 miles
--------------------
Equation:
12t + 16t-24 = 40 miles
----
28t = 64
time = 16/7 = 2hr 17 min
Ans: 8:00 + 2hr 17 min = 10:17AM
-----------
Cheers,
Stan H.
--------------